OpenStudy (anonymous):

Let R represent the region bounded by y = sinx and y = x^4. Find: (a) the area of R; (b) the volume of the solid whose base is R, if all cross sections perpendicular to the x-axis are isosceles triangles with height 3; (c) the volume of the solid formed when R is rotated around the x-axis.

OpenStudy (anonymous):

Let's graph two function to see which the intersection

OpenStudy (anonymous):

they both intersect the x-axis at x=0 and x=1

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\sin(x)dx - \int\limits_{?}^{?}x^4 dx\]

OpenStudy (anonymous):

I am not sure if they intersect at 1 sin[1]=not 1 [1]^4=1 I guess we can use small angel theorem sin[θ]=θ sin[1]=1 (1)^4=1

OpenStudy (anonymous):

oh yeah crap i forgot

OpenStudy (anonymous):

wow this solution is ridiculous

OpenStudy (anonymous):

Yeah, its like .95 not quite 1. Close enough.